GoConqr has a new look! Let us know what you think.

Flashcards

probability

Tyde D GEG 100 English 10 cards

Cards in this deck (10)

  1. 1what is 1/8 as a decimal fraction0.125
  2. 2Which of these numbers cannot be a probability? a) -0.00001 b) 0.5 c) 1.001 d) 0 e) 1 f) 20%a) and c)
    probabilities are between 0 and 1
  3. 3If you toss a coin ten times, how many possible outcomes are there?2^10 = 1024
  4. 4how many coins do you have to toss before the probability of throwing 'all the same' (heads or tails) falls below 5%?the sequence goes
    one toss 1/1 = 1
    two toss 1/2 = 0.5
    three toss 1/4 = 0.25
    four toss 1/8 = 0.125
    five toss 1/16 = 0.062
    six toss 1/32 = 0.031
    answer: 6 coins because 0.031 is smaller than 0.05
    to do this on a calculator use trial and error:
    2^6 = 64
    2/64 = 0.031
  5. 5If you roll four dice, what is the probability of obtaining four numbers that are the same?workings out:
    1) find 6^4 because there are 6 possible outcomes and 4 dice are rolled. answer = 1296
    2) so there are 6 possible correct roles so 6/1296 = 0.0046.
    slightly less than 5 in a thousand
  6. 6true or false: a massive difference in sample data is likely to produce a probability close to 1false:
    a massive difference produces a probability close to 0.
    data that are close to eachother produces a probability close to 1
  7. 7what is the probability of rolling two dice and obtaining two numbers that are the same?36 possible outcomes (6x6)
    there are 6 ways of producing two numbers the same (the six faces of the dice), so the answer is 6 in 36 or 1 in 6
  8. 8In a multiple-choice examination where there are five possible answers for each question, what is the probability of scoring four out of four correct purely by chance?5^4 = 625
    this is LESS THAN 1 in 500
  9. 9you have asked a random sample of ten people the same question and all of them answered yes rather than no. What are the chances of obtaining a ratio of ten to zero simply by chance? About one in ten About one in a hundred Slightly more than one in a thousand Slightly less than one in a thousand Less than one in ten thousandmethod:
    1) 2^10 = 1024
    the answer is slightly LESS than one in a thousand. The chances of something happening one in 1024 times is less than the chances of something happening one in a thousand times.
  10. 10There is clearly an extremely high possibility that the difference is just luck so the p-value must be close to _____?one

Information

Description
GEG 100 Flashcards on probability, created by Tyde D on 11/12/2018.
Created by
Tyde D
Views
0
Rating

Feedback

What is it about?

What do you think of the new look?

How do you like it?